OpenStudy (anonymous):

Given f(x) = the quantity of x plus 4, divided by 6, solve for f-1(7).

OpenStudy (anonymous):

what does this all mean? D:

OpenStudy (anonymous):

\[f(x)=(x+4)/6\] Is this correct

OpenStudy (anonymous):

this is prettier :> \[\Large f(x) = \frac{x+4}{6}\]

OpenStudy (sheraz12345):

f(x)=(x+4)/6 ... Ttake its inverse Its inverse would be i guess 6x-4 insert 7 in 6x-4

OpenStudy (anonymous):

thats correct @PeterPan

OpenStudy (anonymous):

but do you know how to find inverses? :)

OpenStudy (sheraz12345):

@PeterPan are u asking me ?

OpenStudy (anonymous):

@Sheraz12345 No, silly, I'm asking the asker, since you basically made the inverse appear by magic :D

OpenStudy (sheraz12345):

Ohh sorry :p

OpenStudy (anonymous):

1 24 32 38 these are the answers and i got 32

OpenStudy (anonymous):

But how'd you get 32?

OpenStudy (sheraz12345):

Its 38 .. How'd u get 32

OpenStudy (anonymous):

thats how i calculated it. so the anwer is 38

OpenStudy (aum):

@Sheraz12345 You keep supplying answers without helping the person learn.

OpenStudy (sheraz12345):

Well @aum this is the first time i did such a thing .. Wont happen again